OpenStudy (anonymous):

graph the equation and identify the y-concept y=1/4x

OpenStudy (anonymous):

is this like y=(1/4)x ? y=(1/4)x =x/4 if you wanna graph this start maybe with various x=-4,-3,-2,-1,0,1,2,3,4 you will get y=-1,-3/4,-1/4,0, 1/4, 1/2 ,3/4, 1 then plot them and graph them... hope that helps